# An algebraic criterion for the vanishing of bounded cohomology

### v4

@@ 1. Introduction
In a companion paper [CFFLM24], we survey all the displacement techniques available in the literature for the study of bounded cohomology and stable commutator length and we discuss the implications between them. In particular, to our knowledge having commuting cyclic conjugates turns out to be the weakest condition that ensures $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity.
@@ 1.1.1. The stable mapping class group $\Gamma_{\infty}$
On the other hand, Kotschick first proved that the stable commutator length of $\Gamma_{\infty}$ vanishes [Kot08, Theorem 3.1] and then Bowden proved that $H^{2}_{b}(\Gamma_{\infty};\mathbb{R})$ is zero [Bow11, Prop. 4.7]; this together with the fact that stable classes in stable groups tend to unbounded classes led Bowden to conjecture that the bounded cohomology of the stable mapping class group vanishes in all positive degrees:
@@ 1.1.2. Stable braid group and the bounded cohomology of the braided Ptolemy–Thompson group
The stable braid group $B_{\infty}$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. The same is true for its commutator subgroup $[B_{\infty},B_{\infty}]$.
Recall that Thompson’s group $T$ can be identified with the asymptotically rigid mapping classes of the two-dimensional disc $D$ [Kim16]. As an immediate consequence of Corollary 1.6, we can fully compute the bounded cohomology of the braided Ptolemy–Thompson group $T^{\ast}$ [FK08] with trivial real coefficients. Recall that the group $T^{\ast}$ consists in all asymptotically rigid mapping classes of $D^{\ast}$, where $D^{\ast}$ denotes $D$ with a countable infinite locally finite collection of punctures [Kim16].
@@ 1.1.3. Braided Thompson groups
The braided Thompson group $bV$ was introduced independently by Brin [Bri07] and Dehornoy [Deh06], and can be realized as a dense subgroup of the mapping class group of a disc minus a Cantor set [SW21]. Analogously to the braided Ptolemy–Thompson group, there exists an epimorphism $bV\to V$ whose kernel $bP$ is a direct union of pure braid groups. However this directed union is not the standard one, and corresponds to splitting strands, rather than adding unbraided strands. In fact, both $bV$ and $bP$ were shown to admit an infinite-dimensional space of homogeneous quasimorphisms [FFLZ22].
On the other hand, $bV$ contains a subgroup $\widehat{bV}$, also defined by Brin [Bri07], which can be seen as a braided version of a point stabilizer in $V$. The commutator subgroup of $\widehat{bV}$ was shown to have commuting conjugates [FFLZ22], and it turns out that it also has commuting cyclic conjugates. So we obtain:
@@ 1.3. Stability
The previous results have interesting implications for the theory of homological stability (we refer the reader to work of Randal-Williams and Wahl [RWW17] and the references therein). A family of groups $\Gamma_{1}\hookrightarrow\Gamma_{2}\hookrightarrow\cdots$ satisfies homological stability if, for all $i\geq 1$, the maps $H_{i}(\Gamma_{n})\to H_{i}(\Gamma_{n+1})$ are eventually isomorphisms. Homological stability has been an important tool for studying the homology of certain families of groups. Typically the approach splits into three steps: Prove homological stability, then compute the homology of the colimit group $\textup{colim}_{n\to\infty}\Gamma_{n}$, and use it to compute the homology of the groups $\Gamma_{n}$, at least for $n$ large enough.
@@
Theorem 1.3 therefore shows that the colimit groups coming from homologically stable families tend to be $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. This provides an obstacle for adapting homological stability arguments to the setting of bounded cohomology, and answers a question of Kastenholz–Sroka [KS23, Question 1.4] about the scope of this approach, for the homologically stable families above. Let us however remark that bounded cohomological stability has been proved and applied with success in various contexts, where the bounded cohomology of the colimit groups plays no role [Mon04, Mon10, DlCMH23a, DlCMH23b, DlCM23, Kas23, KS23].
@@ 1.4. Transformation groups
The bounded cohomology of transformation groups as in Corollary 1.12 is well-studied in the cases of compact (or more generally finite-volume) manifolds [GG04, Pol06, Py06, EPP12, BM19, Nit21, BM22, CGHM+22, BM23]. In those cases it is generally non vanishing, even in high degree, in stark contrast with our result. Note that the inclusion $\textup{Diff}_{\operatorname{\textup{vol}}}(\mathbb{D}^{n},\partial\mathbb{D}^{n})\to\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n})$ induces an isomorphism in ordinary cohomology [McD83a]. However, this is not the case in bounded cohomology, at least for $n=2$ and degrees $2$ and $3$ [BG92, Kim23] (as far as we know, nothing is known about the bounded cohomology of $\textup{Diff}_{\operatorname{\textup{vol}}}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ for $n\geq 3$ [BIP08, Open Problem, p. 230]).
@@ 1.5.1. Piecewise linear groups of the interval, and piecewise projective groups of the line
These groups have a rich subgroup structure: They both contain the famous Thompson’s group $F$ [CFP96], and the latter contains Monod’s groups $H(A)$ where $A$ is a dense subring of $\mathbb{R}$ [Mon13, Mon23], as well as the Lodha–Moore group $G_{0}$ [LM16]. The groups $H(A)$ where $A$ is a dense subring of $\mathbb{R}$, and $G_{0}$, are examples of non-amenable groups without non-abelian free subgroups (and the Lodha–Moore group $G_{0}$ is a finitely presented and type $F_{\infty}$ example). These groups have the striking property that they do not contain non-abelian free subgroups, yet do not satisfy a law [BS85]. Their homology has been computed in various cases, and it is typically non-zero in infinitely many degrees [BG84, GS87, Gre89, Ste92, Bro06].
@@ 1.5.3. Interval exchange transformations
[…]
@@
[…]
@@ 1.7. Methods
The technical tool underlying these results is the vanishing modulus, a numerical invariant, introduced independently by different teams [MN23, FFLM23], that attaches a quantity to a vanishing bounded cohomology group. In particular, the bounded cohomology vanishes in degree $n$ if and only if its $n$-th vanishing modulus is finite [FFLM23]. The importance of the vanishing modulus is encoded in the following: A direct union of boundedly acyclic groups stays boundedly acyclic if the vanishing moduli of all the groups are uniformly bounded [FFLM23, Proposition 4.13]. Here we generalize this result to $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity. Therefore the technical core of this paper consists in a further study of the vanishing modulus, and in particular in a proof that the vanishing modulus of the class of permutational wreath products $\Gamma\wr_{X}A$ is uniformly bounded. To this end, we first give a complete proof of the $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity of permutational wreath products, with the additional assumption that the action $A\curvearrowright X$ has infinite orbits [Mon22, Section 4.2]. Then, we have to show that those groups are not only $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic (whence with finite vanishing modulus), but that their vanishing modulus is uniformly bounded independently of $\Gamma,X,A$ or even $E$. The last step, which is purely algebraic, consists in expressing groups with commuting cyclic conjugates as a directed union of quotients of groups of the form $\Gamma\wr_{X}A$ by amenable normal subgroups. This requires a careful inductive construction, since $A$ needs to be infinite for the previous results to hold, but the obvious wreath products arising from the definition of commuting cyclic conjugates may well have finite acting group.
@@ Acknowledgements
The authors are indebted to Javier Aramayona, Jon Berrick, Benjamin Bode, Jonathan Bowden, Luigi Caputi, Carlo Collari, Pierre de la Harpe, Rodrigo De Pool, Dominik Francoeur, Camille Horbez, Thorben Kastenholz, Nicolás Matte Bon, Nicolas Monod, Sam Nariman, Pierre Py, Oscar Randal-Williams, George Raptis, Sven Sandfeldt, Arpan Saha, Lauro Silini, Marco Varisco, Xiaolei Wu and Matt Zaremsky for useful conversations.
@@
This work has been supported by the Madrid Government (Comunidad de Madrid - Spain) under the multiannual Agreement with UAM in the line for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). This work has been funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Geometry and topolgy of manifolds”, proposal code 2022NMPLT8 - CUP J53D23003820001.
@@ 2.1. Bounded cohomology with dual coefficients
that sends bounded functions to bounded functions (we will typically simply denote $\delta$ when the degree is clear from the context). By construction $\Gamma$ acts on $C_{b}^{\bullet}(\Gamma;E)$ as follows: For every $f\in C_{b}^{\bullet}(\Gamma;E)$,
@@
    Λ=($\Lambda\wr\mathbb{Z}=\big{(}\bigoplus_{\mathbb{Z}}\Lambda)\rtimes\mathbb{Z}$
@@
    Every countable group embeds into a $\operatorname{\mathfrak{X}^{\{\mathbb{R}\}}}$-boundedly acyclic group, and there exists a finitely presented $\operatorname{\mathfrak{X}^{\{\mathbb{R}\}}}$-boundedly acyclic group that contains all finitely presented groups [FFLM21];
@@ 2.2. Highly ergodic Zimmer-amenable actions
We recall in this section how to compute bounded cohomology using highly ergodic Zimmer-amenable actions. Given a non-singular (i.e. measure class preserving) action of a group $\Gamma$ on a standard Borel measure space $S$, and a Banach $\Gamma$-module $E$, there is an induced action of $\Gamma$ on $L^{\infty}(S,E)$, defined in the same way as the action of $\Gamma$ on $C^{n}_{b}(\Gamma;E)$ (see Equation (1)).
@@
[…]
@@
[…]
@@ 2.4. Direct unions
proving our claim. ∎
@@
Since $B_{H}^{z}\subset B_{H^{\prime}}^{z}$ whenever $H^{\prime}\leq H$, and the system is directed, the finite intersection property will follow as soon as each $B_{H}^{z}$ is non-empty. Now, fix $H\leq\Gamma$ and let $i$ be such that $H\leq\Gamma_{i}$. Since $\|H^{n}_{b}(\Gamma_{i};E)\|\leq K$, there exists $b\in C^{n-1}_{b}(\Gamma_{i},E)^{\Gamma_{i}}$ such that $\delta^{n-1}(b)=z|_{\Gamma_{i}}$ and $\|b\|_{\infty}\leq K\cdot\|z\|_{\infty}$; extending $b$ by 00 we obtain an element in $B_{H}^{z}$, which concludes the proof. ∎
@@
A similar result for trivial real coefficients and unrestricted products was stated by Monod and Nariman [MN23, Proposition 2.6]. Here we provide an elementary, self-contained proof, which allows for coefficients as well. The core of the proposition lies in the following more general fact. Recall that a surjective homomorphism $r\colon\Gamma\to\Lambda$ is a retraction if there exists an injective homomorphism $\iota\colon\Lambda\to\Gamma$ such that $r\circ\iota=\textup{id}_{\Lambda}$.

### v5

@@ Abstract.
Key words and phrases:
Bounded cohomology, bounded acyclicity, commuting cyclic conjugates, wreath products, colimit groups, homological stability, transformation groups, big mapping class groups, Thompson groups, braid groups, linear groups
2020 Mathematics Subject Classification:
Primary: 20J05, 57M07; Secondary: 20E22, 57M60, 57S05
@@ 1. Introduction
In a companion paper [CFFLM25], we survey all the displacement techniques available in the literature for the study of bounded cohomology and stable commutator length and we discuss the implications between them. In particular, to our knowledge having commuting cyclic conjugates turns out to be the weakest condition that ensures $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity.
@@ 1.1.1. The stable mapping class group $\Gamma_{\infty}$
Moreover, Kotschick first proved that the stable commutator length of $\Gamma_{\infty}$ vanishes [Kot08, Theorem 3.1] and then Bowden proved that $H^{2}_{b}(\Gamma_{\infty};\mathbb{R})$ is zero [Bow11, Prop. 4.7]; this together with the fact that stable classes in stable groups tend to unbounded classes led Bowden to conjecture that the bounded cohomology of the stable mapping class group vanishes in all positive degrees:
@@ 1.1.2. Stable braid group and the bounded cohomology of the braided Ptolemy–Thompson group
The stable braid group $B_{\infty}$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. The same is true for its commutator subgroup $B_{\infty},B_{\infty}]$.
Recall that Thompson’s group $T$ can be identified with the asymptotically rigid mapping classes of the two-dimensional closed disc $D$ [Kim16]. As an immediate consequence of Corollary 1.6, we can fully compute the bounded cohomology of the braided Ptolemy–Thompson group $T^{\ast}$ [FK08] with trivial real coefficients. Recall that the group $T^{\ast}$ consists in all asymptotically rigid mapping classes of $D^{\ast}$, where $D^{\ast}$ denotes $D$ with a countable infinite locally finite collection of punctures [Kim16].
@@
where the degree of $x$ is $2$. Under this isomorphism, $x$ corresponds to the Euler class of the defining action on the circle of the quotient $T$.
@@ 1.1.3. Braided Thompson groups
The braided Thompson group $bV$ was introduced independently by Brin [Bri07] and Dehornoy [Deh06], and can be realized as a dense subgroup of the mapping class group of a disc minus a Cantor set [SW21]. Analogously to the braided Ptolemy–Thompson group, there exists an epimorphism $bV\to V$ whose kernel $bP$ is a direct union of pure braid groups. However this directed union is not the standard one, and corresponds to splitting strands, rather than adding unbraided strands. In fact, both $bV$ and $bP$ were shown to admit an infinite-dimensional space of homogeneous quasimorphisms [FFLZ24a].
On the other hand, $bV$ contains a subgroup $\widehat{bV}$, also defined by Brin [Bri07], which can be seen as a braided version of a point stabilizer in $V$. The commutator subgroup of $\widehat{bV}$ was shown to have commuting conjugates [FFLZ24a], and it turns out that it also has commuting cyclic conjugates. So we obtain:
@@ 1.3. Stability
The previous results have interesting implications for the theory of homological stability (we refer the reader to work of Randal-Williams and Wahl [RWW17] and the references therein). A family of groups $\Gamma_{1}\hookrightarrow\Gamma_{2}\hookrightarrow\cdots$ satisfies homological stability if, for all $i\geq 1$, the maps $H_{i}(\Gamma_{n})\to H_{i}(\Gamma_{n+1})$ are eventually isomorphisms (this is usually meant with respect to trivial integral coefficients). Homological stability has been an important tool for studying the homology of certain families of groups. Typically the approach splits into three steps: Prove homological stability, then compute the homology of the colimit group $\textup{colim}_{n\to\infty}\Gamma_{n}$ (which in this setting is a direct limit), and use it to compute the homology of the groups $\Gamma_{n}$, at least for $n$ large enough.
@@
These results therefore show that the colimit groups coming from homologically stable families tend to be $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic (in fact, it was suggested to us by Thorben Kastenholz that Theorem 1.3 might apply every time that the homologically stable system has an underlying $E_{\infty}$ structure, see [RWW17]). This provides an obstacle for adapting homological stability arguments to the setting of bounded cohomology, and answers a question of Kastenholz–Sroka [KS23, Question 1.4] about the scope of this approach, for the homologically stable families above. Let us however remark that bounded cohomological stability has been proved and applied with success in various contexts, where the bounded cohomology of the colimit groups plays no role [Mon04, Mon10, DlCMH25, DlCMH23, DlCM23, Kas25, KS23].
@@ 1.4. Transformation groups
The bounded cohomology of transformation groups as in Corollary 1.12 is well-studied in the cases of compact (or more generally finite-volume) manifolds [GG04, Pol06, Py06, EPP12, BM19, Nit21, BM22, CGHM+22, BM25]. In those cases it is generally non vanishing, even in high degree, in stark contrast with our result. Note that the inclusion $\textup{Diff}_{\operatorname{\textup{vol}}}(\mathbb{D}^{n},\partial\mathbb{D}^{n})\to\operatorname{\textup{Diff}_{c,\operatorname{\textup{vol}}}}(\mathbb{R}^{n})$ induces an isomorphism in ordinary cohomology [McD83a]. However, this is not the case in bounded cohomology, at least for $n=2$ and degrees $2$ and $3$ [BG92, Kim25] (as far as we know, nothing is known about the bounded cohomology of $\textup{Diff}_{\operatorname{\textup{vol}}}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ for $n\geq 3$ [BIP08, Open Problem, p. 230]).
@@ 1.5.1. Piecewise linear groups of the interval, and piecewise projective groups of the line
These groups have a rich subgroup structure: They both contain the famous Thompson’s group $F$ [CFP96], and the latter contains Monod’s groups $H(A)$ where $A$ is a dense subring of $\mathbb{R}$ [Mon13, Mon25], as well as the Lodha–Moore group $G_{0}$ [LM16]. The groups $H(A)$ where $A$ is a dense subring of $\mathbb{R}$, and $G_{0}$, are examples of non-amenable groups without non-abelian free subgroups (and the Lodha–Moore group $G_{0}$ is a finitely presented and type $F_{\infty}$ example). These groups have the striking property that they do not contain non-abelian free subgroups, yet do not satisfy a law [BS85]. Their homology has been computed in various cases, and it is typically non-zero in infinitely many degrees [BG84, GS87, Gre89, Ste92, Bro06].
@@ 1.5.2. Highly transitive actions
where the degree of $x$ is $2$, and $x$ corresponds to the Euler class of the defining action on the circle of $\Gamma$.
@@
Since the first version of this paper, this approach has been used successfully in more contexts, always using Theorem 1.3 as a starting point: see [WWZZ25, FFMN24, Zho24].
@@ 1.5.3. Interval exchange transformations
Another group of interest arising from one-dimensional dynamics is the group of interval exchange transformations $\operatorname{\textup{IET}}([0,1))$. This is defined as the group of right-continuous bijections of $[0,1)$ given by cutting the interval into finitely many pieces and rearranging by translations. The group, as well as many generalizations, has been widely studied [DFG13, JMBMdlS18, DFG20, GL22, CL22]. The two main open problems about it are whether it is amenable (known as de Cornulier’s conjecture [dC14]), or contains non-abelian free subgroups (known as Katok’s conjecture [DFG13]). Therefore it is natural to wonder whether $\operatorname{\textup{IET}}([0,1))$ has vanishing bounded cohomology (Question 5.44): This can be seen as an analogue to Grigorchuk’s question about bounded acyclicity of Thompson’s group $F$ [Gri95], which was answered by Monod [Mon22].
We consider instead the group of interval exchange transformations of the half line interval $0,\infty)$, which is defined the same way: In particular we allow only finitely many intervals to be exchanged. Note that this group is amenable, respectively contains non-abelian free subgroups, if and only if $\operatorname{\textup{IET}}([0,1))$ has the same property (Lemma 5.41).
@@
The group $\operatorname{\textup{IET}}([0,\infty))$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, and so are all of its derived subgroups.
@@ 1.7. Methods
The technical tool underlying these results is the vanishing modulus, a numerical invariant, introduced independently by different teams [MN23, FFLM23], that attaches a quantity to a vanishing bounded cohomology group. In particular, the bounded cohomology vanishes in degree $n$ if and only if its $n$-th vanishing modulus is finite [FFLM23]. The importance of the vanishing modulus is encoded in the following: A direct union of boundedly acyclic groups stays boundedly acyclic if the vanishing moduli of all the groups are uniformly bounded [FFLM23, Proposition 4.13]. Here we generalize this result to $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity. Therefore the technical core of this paper consists in a further study of the vanishing modulus, and in particular in a proof that the vanishing modulus of the class of permutational wreath products $\Gamma\wr_{X}A$ is uniformly bounded. To this end, we first give a complete proof of the $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity of permutational wreath products, with the additional assumption that the action $A\curvearrowright X$ has infinite orbits [Mon22, Section 4.2]. Then, we have to show that those groups are not only $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic (whence with finite vanishing modulus), but that their vanishing modulus is uniformly bounded independently of $\Gamma,X,A$ or even the coefficient module $E$. The last step, which is purely algebraic, consists in expressing groups with commuting cyclic conjugates as a directed union of quotients of groups of the form $\Gamma\wr_{X}A$ by amenable normal subgroups. This requires a careful inductive construction, since $A$ needs to be infinite for the previous results to hold, but the obvious wreath products arising from the definition of commuting cyclic conjugates may well have finite acting group.
@@ Acknowledgements
The authors are indebted to Javier Aramayona, Jon Berrick, Benjamin Bode, Jonathan Bowden, Luigi Caputi, Carlo Collari, Pierre de la Harpe, Rodrigo De Pool, Dominik Francoeur, Camille Horbez, Thorben Kastenholz, Nicolás Matte Bon, Nicolas Monod, Sam Nariman, Pierre Py, Oscar Randal-Williams, George Raptis, Sven Sandfeldt, Arpan Saha, Lauro Silini, Marco Varisco, Xiaolei Wu and Matt Zaremsky for useful conversations. They also thank the anonymous referees for their careful reading and helpful comments.
@@
This work has been supported by the Madrid Government (Comunidad de Madrid - Spain) under the multiannual Agreement with UAM in the line for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). This work has been funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Geometry and topology of manifolds”, proposal code 2022NMPLT8 - CUP J53D23003820001.
@@ 2.1. Bounded cohomology with dual coefficients
that sends bounded functions to bounded functions (we will typically simply denote it by $\delta$ when the degree is clear from the context). By construction $\Gamma$ acts on $C_{b}^{\bullet}(\Gamma;E)$ as follows: For every $f\in C_{b}^{\bullet}(\Gamma;E)$,
@@
    $\Lambda\wr\mathbb{Z}=\big(\bigoplus_{\mathbb{Z}}\Lambda)\rtimes\mathbb{Z}$
@@
    Every countable group embeds into a $\operatorname{\mathfrak{X}^{\{\mathbb{R}\}}}$-boundedly acyclic group, and there exists a finitely presented $\operatorname{\mathfrak{X}^{\{\mathbb{R}\}}}$-boundedly acyclic group that contains all finitely presented groups [FFLM24];
@@ 2.2. Highly ergodic Zimmer-amenable actions
We recall in this section how to compute bounded cohomology using highly ergodic Zimmer-amenable actions. Recall that a Borel space is standard if it is Borel isomorphic to the Borel space associated to a Polish space. Given a non-singular (i.e. measure class preserving) action of a group $\Gamma$ on a standard Borel measure space $S$, and a Banach $\Gamma$-module $E$, there is an induced action of $\Gamma$ on $L^{\infty}(S,E)$, defined in the same way as the action of $\Gamma$ on $C^{n}_{b}(\Gamma;E)$ (see Equation (1)).
@@
Let $\Gamma$ be a group and let $S$ be a standard Borel probability $\Gamma$-space. We say that $S$ is highly ergodic Zimmer-amenable, $\operatorname{\textup{HEZA}}$ for short, if $S$ is a Zimmer-amenable $\Gamma$-space and the $\Gamma$-action on $S$ is highly ergodic.
@@
Let $\Gamma$ be a countable group for which there exists a $\operatorname{\textup{HEZA}}$-space. Then $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
@@ 2.4. Direct unions
Since $\delta^{n-1}(\pi^{n-1}_{*}(b))=z$, this proves our claim. ∎
@@
Since $B_{H}^{z}\subset B_{H^{\prime}}^{z}$ whenever $H^{\prime}\leq H$, and the system is directed, the finite intersection property will follow as soon as each $B_{H}^{z}$ is non-empty. Now, fix $H\leq\Gamma$ and let $i$ be such that $H\leq\Gamma_{i}$. Since $\|H^{n}_{b}(\Gamma_{i};E)\|\leq K$, there exists $b\in C^{n-1}_{b}(\Gamma_{i},E)^{\Gamma_{i}}$ such that $\delta^{n-1}(b)=z|_{\Gamma_{i}}$ and $\|b\|_{\infty}\leq K\cdot\|z\|_{\infty}$; extending $b$ by 0 we obtain an element in $B_{H}^{z}$, which concludes the proof. ∎
@@
A similar result for trivial real coefficients and unrestricted products was stated by Monod and Nariman [MN23, Proposition 2.6]. Here we provide an elementary, self-contained proof, which allows for non-trivial coefficients as well. The core of the proposition lies in the following more general fact. Recall that a surjective homomorphism $r\colon\Gamma\to\Lambda$ is a retraction if there exists an injective homomorphism $\iota\colon\Lambda\to\Gamma$ such that $r\circ\iota=\textup{id}_{\Lambda}$.